The average weight of a newborn baby is 6.4 pounds with standard deviation 1.2 pounds.  If a baby girl weigh is in the 90th percentile how much does she weigh?

multiple choice:

a) 6.4

b) 12.8

c) 7.9
Answer #1

We assume that the weight is normally distributed.

Let X be the r.v denoting the weight of a new born baby.

Given P(X<X_0)=0.9

Now \mu=6.4 \;\&\; \sigma =1.2

Thus,Define \; Z:=\frac{X-\mu}{\sigma} \; such \; that \; Z\sim N(\mu=0,\sigma^2 =1)

P(X<X_0)=P\left (\frac{X-\mu}{\sigma} <\frac{X_0-\mu}{\sigma}\right )= P\left (Z <Z_0\right )=0.9.

Now, So from Normal Distribution table we we get Z_0=1.285

\frac{X_0 -\mu}{\sigma }=1.285

\frac{X_0 -6.4}{1.2 }=1.285

X_0 =1.285(1.2)+6.4=7.936\approx 7.9

So option C is correct